OpenStudy (anonymous):

The resistivity of copper 1.72 x 10^-8ohms -m. A copper wire diameter is .5000 cm. What is the resistence of the wire?

OpenStudy (roadjester):

\[R= \rho \frac l A\]

OpenStudy (anonymous):

ok so how do i find the area?

OpenStudy (roadjester):

\[A= \pi r^2\] you know diameter

OpenStudy (anonymous):

So...A= .0625 pi?? But thats just the are of a cirle? Isn't the copper wire a cylinder

OpenStudy (roadjester):

I assumed you were given length. without it, this problem won't work

OpenStudy (roadjester):

also, don't forget to convert the .5 cm to meters

OpenStudy (anonymous):

Oh sorry I left one vital part of the quatione: the lenght of the cooper wire is 15 m

OpenStudy (roadjester):

there ya go

OpenStudy (roadjester):

think you're good?

OpenStudy (anonymous):

mmmhhh..... i need to know how to find the area

OpenStudy (roadjester):

use the area for a circle. you know the diameter, so convert cm to m, divide by 2, and you can get the area of a circle

OpenStudy (anonymous):

is a copper wire in the shape of a cylinder....if so should i find the area of that piticular cylinder?

OpenStudy (roadjester):

\[\huge R=\rho \frac l {\pi {(\frac d 2)}^2}\] make sure your units work out but you don't want to always assume a wire is cylindrical. if you are given a diameter or radius, it's almost always going to be a cylider, but sometimes you may be given a square wire.

OpenStudy (anonymous):

oh...okay thx....I'll give you points.